Prev Next

SW&IT Meeting Minutes 2019-04-08

Attendees: Andrew, David, Igor, Matthieu, Petr, Stephen, Xianyu, Yang

Outlook

Labor Reporting

Report auto-generated 20190325 06:45, Do not edit this table, use entries in smartsheets.

LR20190325 - Bec, Matthieu 80.0
LR20190325 - Colson, Andrew 80.0
LR20190325 - Edmeades, David 72.0
LR20190325 - Sola, Igor 37.0
LR20190325 - Kubanek, Petr 80.0
LR20190325 - Hooper, Stephen 80.0
LR20190325 - Zhang, Yang 78.0
LR20190325 - Zhang, Xianyu 87.0

News from management

  • ExpressMaintenance > MS Access in-house DB (Pat)
  • Daytime testing, bringing back system for night-time ops concerns
  • Proposal and Time Allocation at LBTO (WPC science highlight from ME)
  • sci-arc will request 2x new VMs for new archive, asked zabbix/nagios monitoring

Previous Action Items

Open topics

  • CODING.md - TCS coding practices. See ~pkubanek/tcs/CODING.md on developing machines for a draft
  • Wired network in rm507 - BR requests to add 2 switches "one next to RM507-2 on the north wall would be useful, perhaps a 3rd on the table in the center"

Team reports

PK

  • deamon mode, final touches on hkserver (new HK for old OAC hardware)
  • TCS socket testing showing C++11 powers
  • OSS Umac module improvemens (uses standard TCS Socket class instead of custom one)
  • finally found VimDoxygenToolkit after suspecting for a long time this must exists
  • CODING.md (based on experience from YZ pull request review)

MB

  • OC, WPC, Management
  • EE's interview x2
  • OVMS+tt stand-up, LN, science archive meetings
  • DMS new telemetry
next
  • DMS new telemetry
  • timeseries db
  • phone screening UK candidate

YZ

  • Closed IT#7610 (MODS script issue) after receiving help from Rick Pogge.
  • TCS work:
    • Finished a branch to include the new command "ApplyWaffleAO" (with changes mainly in IIF and AOS part, the major missing part is the AOSupervisor OpCode for the new command, which I think should be generated on the AOSupervisor part and then imported).
    • Created a pull request for review.
    • Next to work on new instrument authorization.
  • AO-Cravat:
    • Some major reorganization of the web interface.
    • Still having trouble with Chart.js time axis setting (won't work even on very simple independent HTML test page), need to resolve this.
    • Still working on tweaking the web interface.
    • Created a pull request to merge recent changes (not everything is in working order, just to sync the Github repo).
    • Needs a meeting with Al Conrad to clarify several design requirements.

DE

  • L4AO
    • Regular script-based generation of pages working well; no one has broken my logic yet
    • Minor edits
  • lbto.org
    • Stephen expressed the desire to move away from external hosting and towards UA; I concur.
    • Looking at availability of modern templates
      • Easy to work with
      • Cheap or even free (~$50 for a year of access, no monetary obligation after that if we don't need support)
      • Already mobile-aware
  • UITS followup; we are "close" to having someone assigned. My contact has been Christopher Gery.
  • Issue with Jim's video card.
    • Only one port is working now
    • I'm using it as a motivation for him to complete his move to the new, much more capable machine that he's had on his desk for months.
    • If it has further issues, we will replace it
  • Sorted out IssueTrak critical ticket email/SMS notifications with an IT engineer
    • Collecting mobile numbers from interested parties. If anyone is complaining about not getting SMS, verify that they sent me their number.
      • Alerts go to the "LBTO_Critical" group
      • Users who wish SMS alerts will also need the "sms email" field filled out in their user details (number@carriertextemail.com)
    • Natural critical event did send SMS alerts out; got several confirmations.
Next
  • L4AO monitoring and scripts
  • Mockup of example lbto.org site
  • Wiki problems: This morning's look shows poor console responsiveness for no clear reason (RAM is fine, load is fine, no obviously weird processes)
  • Hopefully UITS meeting

SH

  • talked to Leroy, all parts for wireless should be there, checked with mountain management, scheduled car and bedroom
  • update SSL cert -- purchased as "it@lbto.org", reset password for it@lbto.org at Namecheap and put in pdb.
  • PEPSI PDU issues -- locked out Ilya. Cleaned up the access to the PDUs.
  • problems with TCS machines rebooting, mounts Ok, but things mounted on top of other things. Probably should go with a systemd mount instead of the fstab mounts.
next:
  • Wireless cabling and infrastructure (up the mountain Monday and Tuesday nights).
  • shares for Matthieu's request, RPMS to feed them.
  • Wiki keeps dying -- definitely a memory issue. David will investigate, while I am on my little trip up the mountain.

AC

  • OAC 5.2 released - logs show successful EPICS channel writing
  • PMS delivered to mountain on Friday
  • Added FACSUM logic to loop through radar images
next
  • Install PMS software on 32-bit machine (OAC?, JET?) on mountain
  • Add FACSUM logic to display current radar image
  • Try different radar image widget layouts on FACSUM
  • ARGOS ITs

IS

Last week
  • PIT
    • Finish INAF pit work. This took most of the week.
  • LUCI
    • LUCI issuetrak 7473. I met with Barry and we agreed it was a configuration issue and on a possible path forward in case this happens again. Issuetrak has been updated.
    • Work some on entrace window dust detection algorithm for LUCI
    • Flat_K_profile_20190315.pngFlat_K_b4_integration_20190315.pngFlat_K_b4_profile_20190320.png
This week
  • PIT
    • Arizona spreadsheet
  • LUCI
    • Working on entrace window dust detection algorithm for LUCI
  • OT
    • Apply the AO/GUIDE star mix coordinate fix to production
I AttachmentSorted ascending Action Size Date Who Comment
Flat_K_b4_integration_20190315.pngpng Flat_K_b4_integration_20190315.png manage 138 K 08 Apr 2019 - 17:38 IgorSola Luci2 Flat on K filter with dust in entrance window with bias removed.
Flat_K_b4_profile_20190320.pngpng Flat_K_b4_profile_20190320.png manage 174 K 08 Apr 2019 - 17:40 IgorSola Luci2 Flat on K filter with entrance window cleaned with fitting line
Flat_K_profile_20190315.pngpng Flat_K_profile_20190315.png manage 130 K 08 Apr 2019 - 17:39 IgorSola Luci2 Flat on K filter with dust in entrance window profile with fitting line
Topic revision: r12 - 08 Apr 2019, IgorSola
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback